OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3?

OpenStudy (anonymous):

-3 is the slope so -3 = -3*5 + C -3 + 15 = C C = 12 y = -3x + 12

OpenStudy (anonymous):

Y+3X=12 as shown by anaria

OpenStudy (anonymous):

3x – y = 12 3x – y = –12 3x + y = –12 3x + y = 12

OpenStudy (anonymous):

oh thnks